A 760 kg spacecraft has total energy -5.2 ? 1011 J and is in circular orbit about the Sun.

find :

(a) Find its orbital radius. M

(b) Find its speed. km/s
Answer #1

E= -K.E. = 1/2 U.

K.E. = -1/2 * m *v^2
v= sqrt(K.E. * 2/m)
v= sqrt(--5.2x10^11 * 2/760) = 36992.18 m/s (answer)

-K.E. =1/2 U.= -G M m/ r'r = 2K.E.
r= -2 G M m/ 2 E
r= -6.67x10^-11 * 1.98892